UT Sep 02 2022 - Night #1 of Restart



Night Info

  • Sunset 18:44
  • Nautical Twilight Ends 19:39
  • Astro Twilight Ends 20:09
  • Moonset 21:26
  • Astro Twilight Starts 04:30
  • Nautical Twilight Starts 05:00
  • Sunrise 05:55
  • Moon ~33%

Staff

  • OSA: D.G.Huerta
  • ISp: C.Bolyard
  • Mtn Manager: P.Hartley
  • ISA/SO: B.Rothberg, A.Becker, R.T. Gatto
  • Technical Expert: O.Kuhn

Night Plan

Telescope Wakeup Part One

  • With authorized instrument - test presets
    • DONE Track (LBC)
    • Acquire (LUCI or PEPSI)
    • Active (LUCI or PEPSI)

LBC Wakeup

  • If clear at sunset - Skyflats
  • DONE Requires LBCs (one or both) released for night use
  • DONE Collimate WRS/DOFPIA - first collimation (0.25hr), then collimate on each move (until TMS is installed) - High
    • Requires Delta-T < 4 deg between ambient and mirror
    • clouds OK, almost any seeing OK
  • Superfocus sequence -PA=0,180, 90, 270 - high (1 hr)
    • Requires seeing 1.2" or better, some clouds (thicker clouds require brighter stars)
  • DONE Rotator Center Checks, start with PA=0, 180 (track mode), then 90,270 - high (1hr)
    • any clouds, seeing < 1.5"
    • DONE first observation RA20 Dec60
    • DONE 2nd elevation
  • Standard photometric fields - medium - (2hr)
    • must be clear, any seeing
    • minimum two airmasses needed
  • Spectro-photometric standards - low - (1hr)
    • must be clear, any seeing
  • Test Science OB - high - (1.5hr)
    • Any conditions (as long as collimation can be done)
    • DONE Sidereal target (0.75hr - including collimation)
    • Non-Sidereal Target (0.75hr - including collimation)
      • 2020 AN3 - visible all night
        • DONE Pass 1
        • DONE Pass 2 w/o guiding
      • 2018 XG5 - end of night

LUCI AGWs 1&2 Wakeup

  • Reconfigure from LBCs to LUCIs - 0.5hr
  • Requires both LUCIs and AGWs (or at least one) released for imaging on-sky engineering
  • AGW1/2 Alignment - 3hrs total
    • Files needed:
      • ssh -X into TO1 as telescope; then ssh - X as tcs@tcs1
      • /home/telescope/TCS/Configuration/GCS/LUCI_L.cfg or LUCI_R.cfg
      • image files: /lbt/data/share/tcs/lbto.YYYMMDD
      • oac@oac: /lbt/oacontrol/current/etc/oacontrol.conf
    • Pupil alignment check - find a star and center it - need Bent Gregorian (1 hr) - need to do before AGW checkout
    • verify SH spots illuminated - critical - (0.5 hr)
    • verify guiding hotspot center - critical - (1 hr)

If LBCs in the beam at end of night - Skyflats in morning twilight


Closed Dome Work

  • LBC Biases (can be done before opening - take set of 25 biases per LBC)
  • LUCI imaging flats - take z,J,H,K,Ks N375 flats at end of night (or during night if closed)

If cannot open

  • LUCI checkouts if needed
  • Test preset with LBCs (Track)
  • Test preset with LUCI
    • Track
    • Acquire
    • Active (simulated guidestar)
  • Test preset with PEPSI
    • PEPSI Pol Acquire
    • PEPSI Pol Active
    • PEPSI PFU Acquire
    • PEPSI PFU Active
  • Other work
    • AO closed dome - execute tasks otherwise scheduled for daytime

Night Log

GoogleDoc from A.Becker

Closed Dome

  • UT 01:09 - humidity at 97% - D.Thompson running N375 PFC imaging mode data collection
  • Current Allsky is thick clouds
  • AllSkyCurrentImage.JPG
  • UT 01:21 - D.Thompson completed PFC calibrations
  • UT 01:23 - reconfiguring for LBCs
  • UT 01:37 - LBCs powered on
  • UT 01:56 - Telescope is closed after sunset - humidity at 89.5% but clouds remain thick
  • UT 02:09 - humidity 83% - but clouds still rollong overhead
  • UT 02:13 - DG Huerta checking on clouds outside
  • UT 02:20 - test preset - may open soon
  • UT 02:32 - opening enclosure

LBC Checkout

Rotator Center
  • UT 02:43 - sending preset to co-pointing field - telescope is not moving, resending again, now moving
  • UT 02:46 - starting doHybrid
  • UT 02:52 - TSS warning AO Wind - finding a field not in the north
  • UT 02:57 - DOFPIA converged - Winds trending down - so will continue with co-pointing and take at least the 0 and 180 PA scripts
    • DIMM 1.6" - LBC Images 1.4"-1.5"
  • UT 03:01 - sending PA0 ob
  • UT 03:03 - sending PA180.ob
  • UT 03:06 - sending PA90.ob
  • UT 03:08 - sending PA270.ob
  • Completed one elevation (20h60d)
  • UT 03:10 - moving to a 2nd field at a different elevation
  • UT 03:12 - running DOFPIA
  • UT 03:15 - not enough pupils in field (also be thick clouds didn't help)
  • UT 03:16 - moving to an SDSSY field RA2044 Dec+01
  • UT 03:18 - humidity 85%
  • UT 03:18 - starting dohybrid
  • UT 03:22 - DIMM showing 1.3"-1.4" - pupils on LBC images look very fuzzy
    • large negative z4 offset - can't find hole LBCR
    • aborting dohybrid (shift+S), clearing active optics on DX
    • IDL on obs2 appears to have frozen
  • UT 03:30 - rerunning dohybrid, with /askme
    • first LBCR pupil images can't see swingarm or hole, DIMM 1.3", delte-temp (ambient-mirror)
    • 2nd image could see swingarm and hole
  • UT 03:47 - copointing
    • DIMM 1.2", LBC images ~1.5"
  • UT 03:48 - PA0.ob
  • Completed 2nd elevation
Test Science OB (Blue Snowball)
  • Blue Snowball Target - DIMM was trending down, but now back up - too poor seeing to do superfocus - running through OBs to test system
  • UT 03:56 - sending co-pointing OB to do DOFPIA
  • UT 04:06 - co-pointing - seeing ~1.5"
  • UT 04:07 - staring science-like OB
  • UT 04:10 - seeing in LBC images 1.2" Uband, 1.1" z-band
  • UT 04:18 - some frames saturated - but showing 1.2" IQ in images
  • UT 04:26 - IQ 1.28" LBC-B, 1.19" LBC-R
  • UT 04:47 - thought there was a jump - but perhaps collimation is not so good - PSFs look triangular
  • UT 04:49 - DIMM is highly vairable hitting as high as 2"
  • UT 04:50 - script ended
Non-Sideral Testing (2020 AN3)
  • UT 04:51 - sending co-pointing OB for collimation
  • UT 05:01 - copointing OB
  • UT 05:05 - sending NS preset, leaving guiding on - will see how this goes with 4min exposures
  • UT 05:15 - think we see the target on Chip2
  • UT 05:31 - OB complete
  • UT 05:31 - re-sending preset NS but with guiding off (guiding failed on 1 or 2 LBCR exposures in the previous OB)
  • UT 05:47 - sky is rapidly closing up on us
  • DONE

Sideral OB Crescent Nebula

  • MW has vanished, picking a field with a large object and lots of stars -central star is R=7.3 try and eek out some IQ info as long as we can, humidiy at 89%
  • UT 06:03 - sending co-pointing OB for collimation
  • UT 06:07 - closing - just becasue we can't see Jupiter......
  • UT 06:34 - reopening
  • UT 06:45 - sending co-pointing OB for collimation
  • UT 06:54 - co-pointing OB
  • UT 06:57 - sending science-like OB - will monitor IQ
    • first LBCB image is out of focus, blue co-pointing image was in-focus - U-Bseell
    • 2nd LBCB image is OK B-Bessell
    • 3rd LBCB now has better IQ than LBCR - 4.5 pix vs 6.3pix V-Bessell
    • 4th LBCB image guiding cutout looks out of focus - U-Bessell out of focus - first imae was U-Bessell
    • the U-Bessell is always out of focus
  • UT 07:15 - created new OB that cycles through all filters in each LBC
  • Uspec OK, U-Bessell not OK, B-Bessell OK, g-slaon OK, V-Bessell OK, r-sloan OK,
    • IT 8695 - Critical (1st of the semester!)
      • What on earth could be causing this? The z offset in mm is verified in the headers to be correct, AND config files are correct!
      • There isn't a 2nd filter out of focus, so this suggests its NOT a filter swap - does U-Bessell have to be inserted in a specific way in the filter holder?
        • IT #690 - reports out of focus, but IT closed as the offsets are correct in the config
        • these were Uspec and just fuzzier but not almost pupil de-focused like what we saw
      • IT #540 - reports B-Bessell out of focus - solution filter put in holder backwards
      • O.Kuhn checking closed-dome flats taken earlier and noticed the counts are not consistent with various filters - if U-Bessell were offset in the wrong spot, then its a domino effect on all the other filters in the wheel.
      • But...if U-Bessell is in empty slot, then g-sloan and r-sloan would be crossed with U-Bessell ... could the U-Bessell just not have been installed...
  • UT 07:27 - humidity 95.4% - closing
  • UT 08:11 - humidity 99.9%
  • UT 09:29 - humidity 93.9% - <1 deg from dewpoint - still closed
  • UT 10:17 - humidity dropped suddenly under 90% - but is slowly bounching back up
  • UT 10:29 - keep jumping back and forth over the 90% humidity line
  • UT 11:44 - Barry calling it done - end of night


-- BarryRothberg - 29 Aug 2022
Topic revision: r9 - 02 Sep 2022, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback